OpenStudy (anonymous):

John and Alan have have a collection of X baseball cards. john has x/4 cards. what fraction of the cards does Alan have?

OpenStudy (ghazi):

total of X cards are there, wherein X/4 is taken by john therefore fraction that Alan has is \[X-\frac{ X }{ 4 }=\frac{ 4X-X }{ 4 }=\frac{ 3X }{ 4 }\]

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:)